The Bon Appetit Bakery makes 4 1/2 times as much revenue on donuts as muffins. If total sales were $44,000 for May, what dollar amount of each was sold?
Answer by oberobic(2304)

x = donut revenue
y = muffin revenue
x + y = 44000
so
y = 44000 - x
.
x = 4.5*y
.
substituting
x = 4.5(44000-x)
multiply though
x = 198000 - 4.5x
add 4.5x to both sides
5.5x = 198000
divide both sides by 5.5
x = 36000
.
y = 44000 - 36000 = 8000
.
Answer:
donut revenue = $36,000
muffin revenue = $8,000